OpenStudy (anonymous):

Which expresses 2700 : 4000 as a decimal? A. 0.07 B. 0.675 C. 0.338 D. 0.54

OpenStudy (anonymous):

@AzureKais

OpenStudy (anonymous):

That is a ratio problem that also can be represented as 2700/4000 = 27/40 = 0.675

OpenStudy (anonymous):

Yeah, all you have to do is divide 2700 by 400 which is .675